OpenStudy (anonymous):

The ratio of the areas of two similar polygons is 225:36. What is the ratio of their corresponding sides

OpenStudy (callisto):

ratio of their corresponding sides = \[\sqrt{225/36} = 5:2\]

OpenStudy (callisto):

alright, that's by similar figure

OpenStudy (anonymous):

ty

OpenStudy (callisto):

yw

OpenStudy (anonymous):

but how is it 5:2 if the sqrt of 36 is 6 ?

OpenStudy (callisto):

you need to simplify the answer!

OpenStudy (anonymous):

oh ok thanks

OpenStudy (callisto):

\[\sqrt{225} = 15\]

OpenStudy (anonymous):

i see now, i just needed to know for future reference thank you so much